OpenStudy (cutestgirlaround):

What is the slope of the line whose equation is −48=2x−8y ?

OpenStudy (1davey29):

Start by putting the equation in y-intercept form (solving for y). After that, use the formula y=mx+b with m being the slope to see your answer. Ex) -4=2x+6y, 6y=-4-2x, y=-(1/3)x-(2/3). Your slope there would be 1/3.

OpenStudy (cutestgirlaround):

??? I'm still confused

OpenStudy (1davey29):

First, solve for y.

OpenStudy (1davey29):

Once you have that, see which value in your equation would be in the same place as the m in y=mx+b. For example, if you solved for y and had y=5x+2, the 5 is in the m's position. In the equation y=mx+b, m is your slope, so in the equation y=5x+2, 5 would be your slope.